OpenStudy (anonymous):

Guys, I really need some help. I am very confused. I need to know how to add and subtract logarithms. logx66 - logx22 I cannot find a calculator that can do this.

OpenStudy (study312):

is x a multiple sign or variable?

OpenStudy (anonymous):

It's like slightly underneath log

OpenStudy (anonymous):

But I couldn't find a way to make it small

OpenStudy (study312):

? could so x66 is under log?

OpenStudy (study312):

i mean so x66*

OpenStudy (anonymous):

The x is and then it's 66 like logx66 but the x is tiny

OpenStudy (welshfella):

OK that means that x is the base of the logarithm one of the rules of logs is log a - log b = log (a/b) - true when the base of the logs is the same

OpenStudy (anonymous):

so logx would stay the same or would it change...?

OpenStudy (welshfella):

logx stays the same logx 66 - logx 22 = logx 66/22 = logx 3

OpenStudy (welshfella):

if you leave a space between the base and the number it makes it easier to understand.

OpenStudy (anonymous):

Oh okay I sort of get it I think... You divided 66/22 to get the 3 for the answer... and if the base is the same it will stay as logx...

OpenStudy (welshfella):

- though you might be able to write x as a suffix in equation editor

OpenStudy (anonymous):

If it is addition would I multiply the numbers in those spots instead of divide?

OpenStudy (welshfella):

yea

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Okay I have a weird one I think I might know the answer but could you check

OpenStudy (anonymous):

log5.4 - 0.054 = log100=2? or log100=10? @welshfella

OpenStudy (welshfella):

i dont get that one is 5.4 supposed to be the base of the log log on its own is taken as log to the base 10

OpenStudy (welshfella):

log 100 is 2 because 10^2 = 100 the log is the power to which the base is taken